OpenStudy (anonymous):

If (8z − 9)(8z + 9) = az2 − b, what is the value of a? The value of a is ____________.

OpenStudy (anonymous):

Im pretty sure its 8 but i want to be sure

OpenStudy (welshfella):

what did you get for the expansion of (8x - 9)(8z + 9)?

OpenStudy (nincompoop):

it is distributive property the coefficients are multiplied \((ax+b) (ax-c) = (ax \times ax) - (ax \times c) + (ax \times c) - (b\times c)\) if you focused on a, it will be \(a \times a = a^2 \)

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

how you got that

OpenStudy (nincompoop):

I am providing the general idea without answering your own homework.

OpenStudy (princeharryyy):

multiply the left hand side and compare the coefficients according to the power of z

OpenStudy (nincompoop):

the coefficients in the first parenthesis will have to be multiplied to the coefficient in the second parenthesis

OpenStudy (princeharryyy):

@lealealea222

OpenStudy (anonymous):

this question confuses me so bad, i dont understand what your saying, im sorry

OpenStudy (nincompoop):

what is a multiplied by a ?

OpenStudy (anonymous):

2?

OpenStudy (nincompoop):

let us use an actual number \(2 \times 2 = ?\)

OpenStudy (anonymous):

4

OpenStudy (nincompoop):

\(4 \times 4 = ? \)

OpenStudy (anonymous):

WAIT its ^2 im so sorry

OpenStudy (anonymous):

its az^2 − b

OpenStudy (nincompoop):

but you are not worried about z nor b you have to worry about a

OpenStudy (anonymous):

ok ok, 8

OpenStudy (anonymous):

no 16

OpenStudy (nincompoop):

how about \(8 \times 8 = ? \)

OpenStudy (anonymous):

64

OpenStudy (nincompoop):

did you notice that \(2 \times 2 = 2^2\) \(4 \times 4 = 4^2\) \(8 \times 8 = 8^2\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so its 64??

OpenStudy (nincompoop):

yes, because the coefficient are both 8

OpenStudy (nincompoop):

if the coefficients were different then you will multiply them as well

OpenStudy (anonymous):

Thankyou :)

OpenStudy (anonymous):

can you help with one more?

OpenStudy (nincompoop):

maybe post it

OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of f[g(3)]? 39 −3 −7 -1

OpenStudy (anonymous):

its like the other one but i dont know it

OpenStudy (anonymous):

@nincompoop